A rectangular field is 40 m longer than it is wide. The field is subdivided into 3 parts by two fences parallel to the width. A total of 560m of fencing was used. What are the dimensions of the field?
Answer by ikleyn(52786)

.
A rectangular field is 40 m longer than it is wide. The field is subdivided into 3 parts by two fences parallel to the width.
A total of 560m of fencing was used. What are the dimensions of the field?
~~~~~~~~~~~~~~~


Let W be the width (in meters).

Then the length is (W+40) meters.


We have 4 (four) parts of the size W  and two parts of the size (W+40).


Altogether they are 560 m.


So, we write this equation

    4W + 2*(W+40) = 560.


Simplify and solve

    4W + 2W + 80 = 560

      6W         = 560 - 80 = 480.

       W                    = 480/6 = 80  meters   (the width)

       Hence, the length is  80 + 40 = 120 meters.


ANSWER.  The field is  80 by 120 meters.

Solved.